one gm of ice at 0 degree c contracts in volume by 0.091 cm^3 on melting.9 gm of metal is heated to 200 c and dropped into an ice calorimeter.The decrease in volume was found to be 0.182 cm^3.Calculate the specific heat capactity of metal.
1
Expert's answer
2020-01-31T10:44:36-0500
"mc(T_2-T_1)=ML"

"(9)c(200-0)=(1)(3.36\cdot 10^5)"

"c=187\frac{J}{kgK}"